# Prompt Recovery for Image Generation Models: A Comparative Study of Discrete Optimizers

### v1

@@ Abstract
Recovering natural language prompts for image generation models, solely based on the generated images is a difficult discrete optimization problem. In this work, we present the first head-to-head comparison of recent discrete optimization techniques for the problem of prompt inversion. We evaluate Greedy Coordinate Gradients (GCG) (Zou et al., 2023), PEZ (Wen et al., 2023), Random Search, AutoDAN Zhu et al. (2023) and BLIP2’s image captioner (Li et al., 2023) across various evaluation metrics related to the quality of inverted prompts and the quality of the images generated by the inverted prompts. We find that focusing on the CLIP similarity between the inverted prompts and the ground truth image acts as a poor proxy for the similarity between ground truth image and the image generated by the inverted prompts. While the discrete optimizers effectively minimize their objectives, simply using responses from a well-trained captioner often leads to generated images that more closely resemble those produced by the original prompts.
@@ 1 Introduction
Images generated by AI models are flooding the internet and the models and prompts used for generation often feel like components of alchemy. Naturally, the academic community aims to better understand the mechanisms at play in these generative systems, with one crucial area focusing on inverting the generative process by recovering natural language prompts from generated images (Fan et al., 2024). Prompt recovery can help control image generation by providing words and concepts from images so users can combine and manipulate prompts for their own use (Wen et al., 2023; Brade et al., 2023) and help analyze the biases and study the possibly harmful associations built into these models (Chin et al., 2023).
Several recently introduced discrete optimization algorithms can invert the generation process and recover prompts from images. Yet, to date, there is no standardized comparison of these methods for prompt inversion from image generation models. In this work, we use a CLIP guided diffusion model to compare and contrast discrete optimizers suited for prompt inversion. Specifically, we measure how PEZ (Wen et al., 2023), GCG (Zou et al., 2023), AutoDAN Zhu et al. (2023), BLIP2’s image captioner (Li et al., 2023), and random search stack up against one another. Our experiments compare the natural language prompt recovered by each method using several metrics and they are followed by a discussion on the proficiency of each method for recovering prompts for image generation.
@@ 2 Related Work
While prompt inversion is an interesting task in its own right, there are two practical motivations to develop strong methods. First, those interested in better controlling the output of image generation models may want to find a prompt from an image as a starting point for their own prompt engineering. Second, by extracting prompts from images, one might better understand the various associations these text-to-image models have and debug them to avoid explicit content generation, for example (Chin et al., 2023). Wen et al. (2023) list iteratively improving prompts as one of the catalysts in their work on discrete optimization for prompt inversion. In fact, image generation can easily and effectively be controlled by strategically refining the text input (Sohn et al., 2023). Refining these prompts can be difficult though and Wen et al. (2023) propose PEZ to optimize natural language inputs to be close in CLIP (Radford et al., 2021) embedding space to target images (we provide more detail on this optimizer in Section 3.1). In contrast, Mahajan et al. (2024) propose an inversion technique that considers the diffusion process instead of relying on CLIP embeddings.
@@
Our focus is on discrete optimization methods for recovering prompts, but discrete optimization over natural language has several other applications like jailbreaking LLMs (Andriushchenko, 2023; Zou et al., 2023; Zhu et al., 2023) and measuring memorization (Schwarzschild et al., 2024; Kassem et al., 2024). In particular, Zou et al. (2023) propose Greedy Coordinate Gradients as a method to find adversarial prompts for LLMs that break their safety alignment. We experiment with this optimizer for image generation inversion as Zou et al. (2023) compare their optimizer to PEZ in their work, but only with the goal of jailbreaking LLMs.
@@
Direct discrete optimization is not the only method for finding viable prompts, however. Several approaches focus on using black box models to sample prompts. Both Zhang et al. (2024) and He et al. (2024) use pretrained language models to extract prompts for given a output across text generation and image generation tasks respectively. Moreover, as we show in this work, even a simple captioner that has not been finetuned for prompt generation often outperforms discrete optimization methods.
@@
Yet, directly optimizing the prompts for inversion tasks remains an important area of study. Similar to prior work on counterfactual explanations (Verma et al., 2020), directly optimizing inputs for a desired outputs helps to better understand the decision boundaries of classifiers (Ribeiro et al., 2016). While generative models are not classifiers with explicit decisions and accuracy metrics, they are constantly making decisions on their representations based on the prompts. From background color to subject ethnicity, discrete optimization methods may provide a useful understanding of the relationship between prompts and images (Williams et al., 2024).
@@
In this work, we emphasize solidifying ways of comparing discrete optimizers for image generation tasks. Even with the rise of novel discrete optimization methods (Zou et al., 2023; Wen et al., 2023; Andriushchenko, 2023), standard comparison on prompt recovery from image generation is missing. We focus on a holistic benchmark that focuses not only on the similarity between prompt and image, but also on the similarity among images generated by the inverted prompts which to the best of our knowledge has not been standardized for this setting.
3 The Optimizers
@@
To best introduce the optimizers we study, it is critical to pose the prompt inversion problem formally. Consider a tokenizer that maps from natural language to sequences of integer valued tokens corresponding to a list of indices in a vocabulary of tokens $\mathbb{T}$. Let $x\in\mathbb{T}^{s}$ be a length $s$ sequence of tokens. Next, let $\mathbf{E}\in\mathbb{R}^{|\mathbb{T}|\times d}$ be a matrix whose rows are $d$-dimensional embedding vectors, one for each token in the vocabulary. To embed a sequence $x$, we can define $\mathbf{X}\in\{0,1\}^{s\times|\mathbb{T}|}$ s.t. $\mathbf{X}\mathbf{1}_{|\mathbb{T}|}=\mathbf{1}_{s}$ to be a matrix whose rows are one-hot encodings of the integers in the sequence $x$ allowing the product $\mathbf{XE}$ to define an $s\times d$ embedding of $x$.
@@
As the name suggests, prompt inversion techniques seek to find the sequence of tokens $x$, or equivalently their corresponding one-hot encodings $\mathbf{X}$, that solve the inverse equation, $\mathcal{M}^{-1}(Y)$, where $\mathcal{M}$ is a stochastic generative model that maps a sequence of tokens $x$ to an image $Y$. Typically we express the solution as the minimizer of some loss function $\mathcal{L}$, or the solution to the following optimization problem.
@@
$\underset{\mathbf{X}\in\{0,1\}^{s\times|\mathbb{T}|}}{\text{argmin}}\quad$ $\mathcal{L}(\mathcal{M}(\mathbf{XE}),Y)$ (1)
[…]
@@
Moreover, computing the gradient through the full generation model $\mathcal{M}$ may be too expensive (for example when $\mathcal{M}$ is a diffusion model and every forward pass may take multiple seconds), but prior work uses CLIP (Radford et al., 2021) to encode images and text in a shared latent space. Some of the methods we examine operate wholly within CLIP’s latent space to compute the loss between the prompt (text) and the target (image). These methods approximate Equation 1 by solving the following problem where $\mathcal{L}_{\text{CLIP}}$ is a similarity loss defined over CLIP embeddings.
$\underset{\mathbf{X}\in\{0,1\}^{s\times|\mathbb{T}|}}{\text{argmin}}\quad$ $\mathcal{L}_{\text{CLIP}}(\mathbf{XE},Y)$ (2)
[…]
@@ 3.1 PEZ
The first approach we consider is PEZ (Wen et al., 2023), a version of projected gradient descent where descent steps are made in the continuous embedding space. The gradients of the objective in Equation 2 are evaluated at points in embedding space corresponding to real tokens, but the trajectory of the iterates may deviate from the set of hard token embeddings.
More formally, let $\mathtt{Proj}_{\mathbf{E}}(\cdot)$ be an operator that projects vectors (or matrices row-wise) from $\mathbb{R}^{d}$ to their nearest row-vector of $\mathbf{E}$, and let $\xi_{i}\in\mathbb{R}^{s\times d}$ be a soft prompt. As an iterative gradient-based optimizer, PEZ produces a sequence of iterates $[\xi_{0},\xi_{1},...,\xi_{n}]$ as it solves the minimization problem in Equation 2. To update from $\xi_{i}$ to $\xi_{i+1}$, PEZ computes the gradient of the loss at the hard prompt $\mathtt{Proj}_{\mathbf{E}}(\xi_{i})$ and takes a step in the direction of this gradient from the soft prompt $\xi_{i}$. Thus, PEZ gives a fast, lightweight method of discrete optimization while still allowing the use of gradient-based descent to approximately solve the problem in Equation 1. For more information, see Algorithm 1 as described by Wen et al. (2023).
@@ 3.2 Greedy Coordinate Gradients
Greedy Coordinate Gradients (GCG) (Zou et al., 2023) is an alternative method for optimizing over the discrete vocabulary using the gradients of the objective with respect to the matrix $\mathbf{X}$ in Equation 2. In particular, we compute the gradient of the loss with respect to $\mathbf{X}$, which is a matrix of the same shape that approximately ranks token swaps. As each entry in a given row of $\mathbf{X}$ corresponds to a token in the vocabulary, each row $i$ in its gradient relays to us how influential changing the token $x_{i}$ to each other token in the vocabulary might be in lowering the loss. More formally, we compute $\nabla_{\mathbf{X}}\mathcal{L}_{\text{CLIP}}(\mathcal{M}(\mathbf{X}\mathbf{E}),Y)$, then, just as gradient descent methods takes steps in the opposite direction of the gradient, we select a random batch of candidate swaps from the top $k$ largest entries of the negative gradient. A given swap corresponds to a single token change in $x$ and we directly compute the loss for each of these candidates and greedily accept the best one as our new iterate.
@@ 3.3 AutoDAN
AutoDAN (Zhu et al., 2023) is an optimizer proposed as a method of finding human-readable adversarial attacks on aligned language models. The optimizer solves Eq. (1) by iteratively optimizing a single token appended to the current prompt. Given an initial prefix, e.g., “Image of a”, the algorithm searches for the token that follows ‘a’ that minimizes the objective function.
The optimizer incorporates a ‘readability’ objective based on the log probability of the next token given an underlying language model. Similarly to GCG, the algorithm employs a coarse-to-fine search strategy. The algorithm appends an initial token, $\hat{x}$ to the current iterate $x$, and scores each token in the vocabulary according to the following scoring function:
@@
AutoDAN was originally designed for text-to-text language models, where the log probability, $\log(p(x_{i}|x))$ was directly available. However, in this review, we use CLIP to determine the quality of the prompt, which does not inherently compute the log probability. We thus use FUSE Williams and Kolter (2024), a recently proposed approach for solving multi-objective problems across models and tokenizers. FUSE approximates the jacobian of a mapping between the two models. By using the embeddings of a text-to-text language model, such as GPT2 to compute both the log probability, $\log(p(x_{i}|x))$, and the gradient, $\nabla_{\hat{x}_{GPT}}\mathcal{L}_{CLIP}(f([x_{GPT},\hat{x}_{GPT}]))$, where $f$ is the mapping from GPT’s embeddings to CLIP’s embeddings. This allows us to apply a language prior when optimizing a prompt with CLIP. We additionally explore the scenario in which we do not use a language prior, by reverting to the standard case in which we optimize over CLIP embeddings, and fix $p(x_{i}|x)=\frac{1}{|\mathbb{T}|}$.
In our experiments, we run AutoDAN for $16$ steps, which enforces a a maximum token length of $16$ due to one-by-one generation of new tokens. We also utilize a beam search with a beam width of 5.
@@ 3.4 Random Search
In order to address this limitation, rather than randomly sampling from the ball around a given embedding, $\mathbf{XE}$, we simulate the behavior of a random sample, by randomly selecting from an $l_{0}$ ball around each element of the sequence in $\mathbf{XE}$. At every iteration, we randomly select a batch of candidates from all embeddings in the $l_{0}$ ball and greedily accept the best single-token replacement as the next iterate.

### v2

@@ Abstract
Recovering natural language prompts for image generation models, solely based on the generated images is a difficult discrete optimization problem. In this work, we present the first head-to-head comparison of recent discrete optimization techniques for the problem of prompt inversion. Following prior work on prompt inversion, we use CLIP’s Radford et al. (2021) text-image alignment as an inexpensive proxy for the distribution of prompt-image pairs, and compare several discrete optimizers against BLIP2’s image captioner (Li et al., 2023) and PRISM (He et al., 2024) in order to evaluate the quality of discretely optimized prompts across various metrics related to the quality of inverted prompts and the images that they generate. We find that while the discrete optimizers effectively minimize their objectives, CLIP similarity between the inverted prompts and the ground truth image acts as a poor proxy for the distribution of prompt-image pairs – responses from well-trained captioners often lead to generated images that more closely resemble those produced by the original prompts. This finding highlights the need for further investigation into inexpensive methods of modeling the relationship between the prompts for generative models and their output space.
@@ 1 Introduction
Images generated by AI models are flooding the internet and the models and prompts used for generation often feel like components of alchemy. Naturally, the academic community aims to better understand the mechanisms at play in these systems, with one crucial focus on inverting the generative process by recovering prompts from images (Fan et al., 2024).
As the name suggests, given only the output image from a generative model, prompt inversion methods seek to find the prompt that generated the image. While the goal is clear, this problem leads to a wide variety of approaches. For example, given an initial prompt estimate, by strategically refining the text input Sohn et al. (2023) have found that the image generation process can be effectively controlled. Yet, this refining process can be difficult, leading to Wen et al. (2023) leveraging CLIP’s (Radford et al., 2021) embedding space to directly optimize natural language inputs to be close to target images. Similarly, Mahajan et al. (2024) have proposed an inversion technique that backpropagates through intermediate steps of the diffusion process instead of relying on CLIP embeddings In contrast, it has been found that training a captioner on a dataset of prompt-image pairs effectively learns the prompt distribution well enough to act as an inverter.
While prompt inversion is an interesting task in its own right, there are two practical motivations to develop strong methods. First, those interested in better controlling the output of image generation models may want to find a prompt from an image as a starting point for their own prompt engineering. Second, by extracting prompts from images, one might better understand the various associations these text-to-image models have and debug them to avoid explicit content. Yet, to date, there is no standardized comparison of these methods for prompt inversion from image generation models.
In this work, we benchmark several approaches to prompt inversion and attempt to shed light on three primary questions: 1) How closely do the distribution of images generated by each approach align with the distribution generated by the original prompt? 2) Do text-image alignment models like CLIP successfully act as a proxy for the joint distribution of prompt-image pairs? 3) Does the ability for discrete optimization methods to search the input space allow them to outperform the ingrained knowledge of learned models? We investigate these questions by comparing four generic methods of discrete optimization within the prompt space (Zou et al., 2023; Wen et al., 2023; Zhu et al., 2023; Andriushchenko, 2023) against two methods that directly sample from the space (Li et al., 2023; He et al., 2024) and evaluate each across several metrics. We follow our analysis with a discussion on the efficacy of different approaches to approximating and searching within prompt spaces.
@@ 2 Related Work
Here, we focus on discrete optimization methods for recovering image prompts, but discrete optimization over natural language has several other applications including jailbreaking LLMs (Andriushchenko, 2023; Zou et al., 2023; Zhu et al., 2023) and measuring memorization (Schwarzschild et al., 2024; Kassem et al., 2024). In particular, Zou et al. (2023) propose a method to find adversarial prompts for LLMs that break their safety alignment. We experiment with this optimizer for image generation inversion as Zou et al. (2023) compare their optimizer to PEZ in their work, but only with the goal of jailbreaking LLMs.
Whereas prompt optimization strategies in the text generation space have specific goals, such as generating targeted strings, the image generation space has struggled with tractable options for aligning prompts and generated images. While we follow prior work in using CLIP as a proxy model, this choice is primarily driven by the practical challenges of directly optimizing prompts through backpropagation in the diffusion process. The computational requirements for coarse-grained exploration and fine-grained search that discrete optimization often calls for (Parker and Rardin, 2014) would entail generating multiple full images for every candidate prompt att every step. Following the search parameters recommended by (Zou et al., 2023), optimizing a single prompt would require generating a minimum of 512 images per step, which quickly becomes prohibitive over numerous iterations.
Mahajan et al. (2024) have attempted to address this burden by focusing on the similarity between predicted noise residuals at specific diffusion timesteps, rather than generating full images. However, in alignment with prior work on noise inversion (Song et al., 2020; Mokady et al., 2023) the authors find that prompts only have strong influence on the generated image during a narrow range of timesteps. At early timesteps, the image becomes largely “locked in,” so even substantial changes to the prompt have little effect. In contrast, at later timesteps, the stochasticity of the diffusion process leads to large variations in the final image, even when the correct prompt is used. This unpredictability makes it difficult to rely on noise residual comparisons for consistent prompt inversion. Thus, prompt inversion methods (Wen et al., 2023; Williams and Kolter, 2024) often rely on deterministic proxy models like CLIP, which offer a more stable and efficient alternative. Through CLIP’s text-image alignment, we can more reliably approximate the prompt-image distribution without having to address the risk of stochasticity producing divergent results.
@@
Importantly, direct discrete optimization is not the only method for finding viable prompts. Several approaches focus on using black box models to sample prompts. Both Zhang et al. (2024) and He et al. (2024) use pretrained language models to extract prompts for given a output across text generation and image generation tasks respectively. Moreover, as we show in this work, even a simple captioner that has not been finetuned for prompt generation often outperforms discrete optimization methods. In fact, Reade et al. (2023) have found that a captioner fine-tuned on pairs of prompts and the images that they generate can effectively sample prompts that are exceptionally similar to the ground truth.
@@
Despite this performance, we focus on the discrete case as direct discrete optimization can be beneficial for better understanding the behavior of the image generation models. Similarly to prior work on counterfactual explanations (Verma et al., 2020), directly optimizing inputs for a desired outputs helps to better understand the decision boundaries of classifiers (Ribeiro et al., 2016). While generative models are not classifiers with explicit decisions and accuracy metrics, they are constantly making decisions on their representations based on the prompts. From background color to subject ethnicity, discrete optimization methods may provide a useful understanding of the relationship between prompts and images (Williams et al., 2024).
@@
We emphasize solidifying ways of comparing discrete optimizers for image generation tasks. Even with the rise of novel discrete optimization methods, standard prompt recovery comparisons over images are missing. We focus on a holistic benchmark on not only the similarity between prompt and image, but also the similarity among images generated by the inverted prompts which to the best of our knowledge has not been standardized in this setting.
3 Selected Algorithms
@@
To best introduce the optimizers we study, it is critical to pose the prompt inversion problem formally. Consider a tokenizer that maps from natural language to sequences of integer valued tokens corresponding to a list of indices in a vocabulary of tokens $\mathbb{T}$. Let $x\in\mathbb{T}^{s}$ be a length $s$ sequence of tokens. Next, let $\mathbf{E}\in\mathbb{R}^{|\mathbb{T}|\times d}$ be a matrix whose rows are $d$-dimensional embedding vectors, one for each token in the vocabulary. To embed a sequence $x$, we can define $\mathbf{X}\in\{0,1\}^{s\times|\mathbb{T}|}$ s.t. $\sum_{i=1}^{T}\mathbf{X}_{j,i}=1~{}\forall j\in\{1,...,M\}$ to be a matrix of one-hot encoded rows for the integers in the sequence $x$. The product $\mathbf{XE}$ defines an $s\times d$ embedding of $x$.
@@
Prompt inversion techniques seek to find the sequence of tokens $x$, or equivalently their corresponding one-hot encodings $\mathbf{X}$, that solve $\mathcal{M}^{-1}(Y)$, where $\mathcal{M}$ is a stochastic generative model that maps a sequence of tokens $x$ to an image $Y$. Typically we express the solution as the minimizer of some loss function $\mathcal{L}$, or the solution to the following optimization problem.
@@
$\underset{\mathbf{X}\in\{0,1\}^{s\times|\mathbb{T}|}}{\text{argmin}}\quad$ $\mathcal{L}(\mathcal{M}(\mathbf{XE}),Y)\quad\text{s.t.}\quad\begin{aligned} &\sum_{i=1}^{|\mathbb{T}|}X_{j,i}=\textbf{1}~{}\forall~{}j\in\{1,...,s\}\end{aligned}$ (1)
@@
Moreover, computing the gradient through the full generation model $\mathcal{M}$ may be too expensive (for example when $\mathcal{M}$ is a diffusion model forward passes may take multiple seconds), but as emphasized above, prior work often uses CLIP (Radford et al., 2021) to encode images and text in a shared latent space. Some of the methods we examine operate wholly within CLIP’s latent space to compute the loss between the prompt and the target image. These methods approximate Equation 1 by solving the following problem where $\mathcal{L}_{\text{CLIP}}$ is a similarity loss defined over CLIP embeddings.
$\underset{\mathbf{X}\in\{0,1\}^{s\times|\mathbb{T}|}}{\text{argmin}}\quad\mathcal{L}_{\text{CLIP}}(\mathbf{XE},Y)\quad\text{s.t.}\quad\begin{aligned} &\sum_{i=1}^{|\mathbb{T}|}X_{j,i}=\textbf{1}~{}\forall~{}j\in\{1,...,s\}\end{aligned}$ (2)
@@ 3.1 PEZ
The first approach we consider is PEZ (Wen et al., 2023), a version of projected gradient descent where descent steps are made in the continuous embedding space. The gradients of the objective in Equation 2 are evaluated at points in embedding space corresponding to real tokens, but the trajectory of the iterates may deviate from the discrete token set.
More formally, let $\mathtt{Proj}_{\mathbf{E}}(\cdot)$ be an operator that projects vectors (or matrices row-wise) from $\mathbb{R}^{d}$ to their nearest row-vector of $\mathbf{E}$, and let $\xi_{i}\in\mathbb{R}^{s\times d}$ be a soft prompt. As an iterative gradient-based optimizer, PEZ produces a sequence of iterates $[\xi_{0},\xi_{1},...,\xi_{n}]$ as it solves the minimization problem in Equation 2. To update from $\xi_{i}$ to $\xi_{i+1}$, PEZ computes the gradient of the loss at the hard prompt $\mathtt{Proj}_{\mathbf{E}}(\xi_{i})$ and takes a step in the direction of this gradient from the soft prompt $\xi_{i}$ and then calls $\mathtt{Proj}_{\mathbf{E}}(\xi_{i})$ to project back to the space of hard prompts. Thus, PEZ gives a fast, lightweight method of discrete optimization while still using gradient-based descent to approximately solve the problem in Equation 1. For more information, see Algorithm 1 as described by Wen et al. (2023). For a single image, we run PEZ over the CLIP loss for 3000 steps and return the prompt the maximizes the CLIP similarity between the image embedding and the text embedding of the prompt.
@@ 3.2 Greedy Coordinate Gradients
Greedy Coordinate Gradients (GCG) (Zou et al., 2023) is an alternative method for optimizing over the discrete vocabulary using the gradients of the objective with respect to the matrix $\mathbf{X}$ in Equation 2. In particular, we compute the gradient of the loss with respect to $\mathbf{X}$, which is a matrix of the same shape that approximately ranks token swaps. As each entry in a given row of $\mathbf{X}$ corresponds to a token in the vocabulary, each row $i$ in its gradient relays to us how influential changing the token $x_{i}$ to each other token in the vocabulary might be in lowering the loss. More formally, we compute $\nabla_{\mathbf{X}}\mathcal{L}_{\text{CLIP}}(\mathcal{M}(\mathbf{X}\mathbf{E}),Y)$, then, just as gradient descent methods takes steps in the opposite direction of the gradient, we select a random batch of candidate swaps from the top $k$ largest entries of the negative gradient. A given swap corresponds to a single token change in $x$ and we directly compute the loss for each of these candidates and greedily accept the best one as our new iterate. As done with PEZ, we run GCG over the CLIP loss for 3000 steps, returning the best prompt as determined by CLIP similarity between the image embedding and the prompt’s embedding.
@@ 3.3 AutoDAN
AutoDAN (Zhu et al., 2023) was proposed as a method of finding human-readable adversarial attacks on aligned language models. The optimizer solves Eq. (1) by iteratively optimizing a single token appended to the current prompt. Given an initial prefix, e.g., “Image of a”, the algorithm searches for the token that follows ‘a’ that minimizes the objective function. The optimizer incorporates a ‘readability’ objective based on the log probability of the next token given an underlying language model. Similarly to GCG, AutoDAN employs a coarse-to-fine search strategy by appending an initial token, $\hat{x}$ to the current iterate $x$, and scores each token in the vocabulary according to the following scoring function:
@@
AutoDAN was originally designed for text-to-text language models, where the log probability, $\log(p(x_{i}|x))$ was directly available. However, in this review, we use CLIP to determine the quality of the prompt, which does not inherently compute the log probability. We thus use FUSE (Williams and Kolter, 2024), a recently proposed approach for solving multi-objective problems across models and tokenizers. FUSE approximates the jacobian of a mapping between the two models and uses the embeddings of a text-to-text language model, such as GPT2 to compute both the log probability, $\log(p(x_{i}|x))$, and the gradient, $\nabla_{x_{GPT}}\mathcal{L}_{CLIP}(f(x_{GPT}))$, where $f$ maps from GPT’s embeddings to CLIP’s embeddings. This allows us to apply a language prior when optimizing a prompt with CLIP. We additionally explore the scenario in which we do not use a language prior, by reverting to the standard case in which we fix $p(x_{i}|x)=\frac{1}{|\mathbb{T}|}$. In our experiments, we run AutoDAN for $16$ steps, which enforces a a maximum token length of $16$ due to one-by-one generation of new tokens. We also utilize a beam search with a beam width of 5.
@@ 3.4 Random Search
In order to address this limitation, we randomly sample from new tokens from the $l_{0}$ ball around each element in the sequence $\mathbf{XE}$. At every iteration, we select a batch of candidatesand greedily accept the best single-token replacement as the next iterate. We compare the prompt found by Random Searching over the same number of steps as done for PEZ and GCG, determining the best prompt by CLIP similarity in the same way.
3.5 PRISM
PRISM, proposed by He et al. (2024), highlights that text-to-image generation is not a one-to-one mapping – multiple prompts can describe the same image, and many images can correspond to the same prompt. Rather than relying on discrete token space optimization, PRISM optimizes over a distribution of prompts. Inspired by LLM jailbreaking methods (eg. Chao et al., 2023), PRISM leverages in-context learning in vision-language models (VLMs) to iteratively refine the prompt distribution. This process incorporates the history of reference images, generated prompts, output images from an anchor text-to-image model, and evaluations from a VLM judge, using techniques similar to chain-of-thought (Wei et al., 2022) and textual gradient (Pryzant et al., 2023). After $K$ iterations across $N$ parallel streams, the best-performing prompt is selected using the same VLM judge. In our experiments, we use GPT-4-o-mini as the VLM and Stable Diffusion XL-Turbo (Sauer et al., 2023) as the anchor text-to-image model, following He et al. (2024)’s setup with $N=6$ and $K=5$. To ensure fair comparisons, we limit the generated prompts to 20 tokens.